# Targeted Immunisation Thresholds for the Contact Process on Power-Law Trees

### v1

@@ Abstract
The popular scale-free configuration model of a social network has a local limit primarily a Galton-Watson tree with some power law offspring distribution $X$, having tail exponent parameter $\tau>2$ such that $\mathbb{P}(X>x)\sim Ax^{2-\tau}$. It is known that contact process epidemics can propagate on these trees and therefore these networks for arbitrarily small infection rates, and uniformly at random immunising a small positive proportion of vertices will have no effect on this result. So, to reduce the possibility of survival of the contact process on these trees we instead remove (immunise) those with largest degree.
For small $\lambda$ we find sharp thresholds in the maximum permissible degree for the survival probability through three distinct $\tau$ regimes, which by duality would correspond to a transition in the metastable density of the associated network after this targeted removal. Above a threshold, we recover the survival probabilities of [mountford2013metastable] and thus the epidemic with immunisation is similar to without, but below a threshold on the same order the survival probability is severely reduced or zero.
@@
John Fernley [^1] HUN-REN Rényi Alfréd Matematikai Kutatóintézet, Reáltanoda utca 13-15, 1053 Budapest, Hungary. fernley@renyi.hu  Emmanuel Jacob222École Normale Supérieure de Lyon, 46 allée d’Italie, 69007 Lyon, France. emmanuel.jacob@ens-lyon.fr
December 7, 2023
@@ 1 Introduction
On a locally finite graph $G=(V,E)$ modelling individuals and their interactions, the contact process is a Markov process in continuous time $(\xi_{t})_{t\geq 0}$ on $\{0,1\}^{V}$ with $\xi_{t}(v)=1$ (resp. 0) if the individual $x$ is infected (resp. healthy) at time $t$. The dynamics are determined by each infected vertex recovering at rate $1$ and each edge $x\sim y$ with for example $x$ infected and $y$ healthy transmitting the infection (to $y$) at fixed rate $\lambda>0$. On an infinite rooted graph, we will typically start the infection from only the root infected and look at the survival probability of the infection $P_{\rm surv}(\lambda):=\operatorname{\mathbb{P}}(\forall t\geq 0,\xi_{t}\neq\mathbf{0})$.
Consider now a distribution $q$ on $\mathbb{N}$ which:
  1. (i)
    follows a power law, or more precisely $q([k,+\infty))\sim c_{q}k^{2-\tau}$ for some $c_{q}>0$ and exponent $\tau>2$.
  2. (ii)
    has mean larger than one: $\sum kq(k)>1$.
The Galton-Watson tree with offspring distribution $q$ is then supercritical by the second condition (note that when $\tau\leq 3$ we have $\sum kq(k)=+\infty$ so the second condition is always satisfied). It is known that on such trees, the survival probability $P_{\rm surv}(\lambda)$ is positive whatever the positive infection rate $\lambda>0$. This of course is made possible by the presence of high-degree vertices in the Galton-Watson tree, as it is easily seen that the contact process a.s. dies out in finite time on a $d$-regular tree when $\lambda\leq 1/d$.
@@
When $\lambda\to 0$, we can even provide the following precise asymptotics for $P_{\rm surv}(\lambda)$ as an adaptation of Proposition 1.4 in [mountford2013metastable]:
@@
$P_{\rm surv}(\lambda)=\begin{cases}\Theta(\lambda^{\frac{\tau-2}{3-\tau}}),&\quad\text{when }\tau\leq 2.5,\\ \Theta(\lambda^{\frac{2\tau-4}{\log^{\tau-2}(1/\lambda)}}),&\quad\text{when }2.5<\tau\leq 3,\\ \Theta(\lambda^{\frac{2\tau-4}{\log^{2\tau-4}(1/\lambda)}}),&\quad\text{when }\tau>3,\end{cases}$ (1)
where $\Theta$ stands for the usual Landau notation. We now truncate the degree distribution at some level $\mathbbm{t}=\mathbbm{t}(\lambda)$ and define the truncated distribution $\tilde{q}=\tilde{q}(\mathbbm{t})$ on $\{0,\ldots,\mathbbm{t}\}$ as
$\tilde{q}(k)=\begin{cases}q(k),&k\in\{1,\ldots,\mathbbm{t}\}\\ q([\mathbbm{t}+1,\infty)),&k=0.\end{cases}$
This is also the distribution of $X1_{X\leq\mathbbm{t}}$ if $X\sim q$. We are interested in seeing whether this truncation can imply extinction of the infection for small $\lambda$, or at least change the asymptotics of $P_{\rm surv}(\lambda)$ as $\lambda\to 0$, depending on the choice of the truncation level $\mathbbm{t}(\lambda)$. Note that we immediately obtain extinction of the contact process when $\mathbbm{t}\leq 1/\lambda$, and so we will typically be interested in the regime $\mathbbm{t}=\mathbbm{t}(\lambda)>1/\lambda$.
@@
Consider a truncation threshold $\mathbbm{t}=\mathbbm{t}(\lambda)\to+\infty$ as $\lambda\to 0$. Write $P^{\mathbbm{t}}_{\rm surv}(\lambda)$ for the survival probability of the contact process with infection parameter $\lambda>0$, starting from only the root infected and run on the Galton-Watson tree with truncated offspring distribution $\tilde{q}(\mathbbm{t}(\lambda))$.
@@
    $P^{\mathbbm{t}}_{\rm surv}(\lambda)=\begin{cases}O\left(e^{-\lambda^{-2+\delta}}\right),&\mathbbm{t}\leq\rho_{1}\lambda^{-2}\log\frac{1}{\lambda},\\ \Theta\left(\frac{\lambda^{2\tau-4}}{\log^{\tau-2}\left(\frac{1}{\lambda}\right)}\right),&\mathbbm{t}\geq\rho_{2}\lambda^{-2}\log\frac{1}{\lambda}.\end{cases}$
@@
    $P^{\mathbbm{t}}_{\rm surv}(\lambda)=\begin{cases}O\left(\lambda^{C}\right),&\mathbbm{t}\leq\rho_{1}\lambda^{-2}\log^{2}\frac{1}{\lambda},\\ \Theta\left(\frac{\lambda^{2\tau-4}}{\log^{2\tau-4}\left(\frac{1}{\lambda}\right)}\right),&\mathbbm{t}\geq\rho_{2}\lambda^{-2}\log^{2}\frac{1}{\lambda}.\end{cases}$
@@
These are asymptotic results as $\lambda\to 0$ and the truncation threshold has $\mathbbm{t}=\mathbbm{t}(\lambda)\to+\infty$. In particular, the truncated degree distribution $\tilde{q}$ will have first moment larger than 1 and the truncated Galton-Watson tree will be supercritical, as is mandatory for the contact process to have any chance to survive. More generally and even when not explicitly stated, many of our results require the hypothesis that $\lambda>0$ is sufficiently small to be satisfied.
@@
In the case $\tau<2.5$, if we also choose $\mathbbm{t}(\lambda)=\rho\lambda^{-\frac{1}{3-\tau}}$, we thus obtain a phase transition in $\rho$ where:
  • For $\rho$ smaller than the critical value, the contact process gets extinct almost surely
  • For $\rho$ larger than the critical value, the truncation does not have any sensible effect on the survival probability, which is at most changed by a multiplicative constant when $\lambda\to 0$ as compared to $P_{\rm surv}(\lambda)$.
@@
In the cases where $\tau\geq 2.5$, we believe a similar phase transition occurs, from almost sure extinction of the contact process, to a survival probability which is essentially the same as without immunisation, and in line with (1). However, our results are not as complete in those cases. First, we do not reveal a single critical value for $\rho$, but rather have to suppose $\rho$ large to get our large $\mathbbm{t}$ estimates, and $\rho$ small to obtain our small $\mathbbm{t}$ upper bounds. Second, if $\tau>2.5$, we cannot even prove extinction of the process for small $\rho$, and instead we only get that the survival probability is significantly decreased. This is due to the difficulty of proving extinction of the contact process under some very unlikely but still possible events, like the truncated tree having many vertices with maximal degree adjacent to the root. Using our current techniques, we still can prove almost sure extinction of the infection, but only if we truncate at a much smaller degree $\mathbbm{t}(\lambda)=O(\lambda^{-2})$.
@@
In the physics literature, these questions have been considered by [pastor2002epidemic, pastor2002immunization] using mean-field heuristics. However, these heuristics struggle to accurately account for the effect of vertices of large degree so targeted immunisation of the large degree vertices is where they are most difficult to justify. This has led to some claims which we are clearly false, for example [pastor2002immunization, Equation 28], and which our results take some steps towards correcting.
@@
Consider immunization instead of truncation, and write $\widetilde{P}^{\mathbbm{t}}_{\rm surv}(\lambda)$ for the survival probability of the contact process with infection parameter $\lambda>0$, starting from only the root infected, on the Galton-Watson tree with offspring distribution $q$, but where every vertex with offspring [^3] We could also immunize every vertex with degree larger than $\mathbbm{t}+1$, instead of every vertex with offspring larger than $\mathbbm{t}$. This would make a difference only for the root (if it has degree exactly $\mathbbm{t}+1$), but would not change our results. larger than $\mathbbm{t}$ is immunized. By definition the immunized vertices cannot get infected or transmit the infection, so these vertices and their offspring can just be removed from the tree. Note this is different from truncation, where only the offspring of the truncated vertices were removed. In principle the contact process could have a different behaviour on the truncated tree, as a truncated vertex can still become infected and reinfect its parent. However we now state that truncation or immunization make no difference when looking at our results.
Proposition.
@@
Let $X$ follow the common offspring distribution $q$. As discussed previously, the truncated tree is simply a Galton-Watson tree with offspring distribution $X1_{X\leq\mathbbm{t}}$. The tree obtained after removal of the immunized vertices is also a Galton-Watson tree, but with a more complicated offspring distribution, which is the main reason for us to consider truncation afterwards. Indeed, let $Y$ have the law of $X$ conditionally on $X<\mathbbm{t}$. Then each vertex has probability $p=\operatorname{\mathbb{P}}(X\leq\mathbbm{t})$ to be a small vertex (with offspring $\leq\mathbbm{t}$), and conditionally on being a small vertex, has offspring distribution $Y$, each child being again a small vertex with probability $p$. Hence, if we condition on the root to be a small vertex, then the tree obtained after removal of the big vertices is a Galton-Watson tree with mixed binomial offspring distribution $\operatorname{Bin}\left(Y,p\right).$
@@ 2.1 Local survival by reinfection from your neighbours, and survival strategies
On the star graph containing a central vertex and $k=k(\lambda)$ leaves, with initially only the center infected, the center will be permanently recovered typically not before time
@@
This simple result dates back to [pemantle1992contact, Theorem 4.1] and has been rewritten or detailed several times since, see e.g. [berger2005starsurvival, Lemma 5.2] or [nguyen2022subcritical, Lemma 3.4]. See also [huangdurrett2018PeriodicTrees, Lemma 4] for a similar upper bound for the expected extinction time of the epidemics. We will use a verion of this result stated for the total time the center is infected, see Lemma A.1, which is a actually a refinement of [nguyen2022subcritical, Lemma 3.4] of independent interest.
On a larger graph, $T(\lambda,\deg(x))$ can still be interpreted heuristically as the time that the infection can survive “locally” at a vertex $x$, using only direct reinfections from the neighbours of $x$. On the truncated tree, the maximal such value is of course $T(\lambda,\mathbbm{t})$, which differs a lot in the different regimes considered in our main theorem:
@@
    Case $\tau<2.5$: Consider $\mathbbm{t}=\rho\lambda^{-\frac{1}{3-\tau}}$ for some constant $\rho>0$. In that case we have $T(\lambda,\mathbbm{t})\to 1$ as $\lambda\to 0$, which is an indication that the reinfections barely have any effect on the spread of the epidemics. We thus expect to survive as in an SIR process, namely if and only if $\lambda\operatorname{\mathbb{E}}[D]>1$.
@@
    Case $\tau=3$: Consider $\mathbbm{t}=\rho\lambda^{-2}$. In that case we have $T(\lambda,\mathbbm{t})=e^{\Theta(\rho)}$ of constant order, so we expect survival if $\lambda\operatorname{\mathbb{E}}[D]$ is bounded below and $\rho$ large enough.
@@
    Case $2.5<\tau\leq 3$. Consider $\mathbbm{t}=\rho\lambda^{-2}\log{\frac{1}{\lambda}}$. We now have $T(\lambda,\mathbbm{t})=\lambda^{-\Theta(\rho)}$. When $\rho$ is large, this time is larger than $\lambda^{-1}$ and thus sufficient to typically infect any given neighbour. We then ensure survival of the infection by observing that in that case the high-degree vertices (with degree of order $\mathbbm{t}$) always form a supercritical Galton-Watson process on which the infection can spread. On the contrary, when $\rho$ is small, we infect any given neighbour with probability only $\lambda^{1-\Theta(\rho)}$, which is not much larger than $\lambda$. In this phase we always have $\lambda\operatorname{\mathbb{E}}[D]\to 0$ as well as $\lambda^{1-\Theta(\rho)}\operatorname{\mathbb{E}}[D]\to 0$ when $\rho$ is small. Therefore the strategy of maintaining the infection at each vertex $x$ for time $T(\lambda,\deg(x))$ before passing the infection to its children will not succeed, which is an indication that the contact process might die out.
  4. (iv)
@@
    $T(\lambda,\mathbbm{t})=\lambda^{-\Theta(\rho)\log(1/\lambda)},$
    a time which is sufficient to infect not only the direct neighbours, but also all nearby vertices up to distance $\Theta(\rho\log(1/\lambda))$. It is thus natural to consider that two high-degree vertices are “connected” if they are separated by less than $\Theta(\rho\log(1/\lambda))$ small-degree vertices, hence obtaining a new “Galton-Watson tree of high degree vertices”. As $\tau>3$ the high-degree vertices are more rare in the original tree, and actually typically at distance $\Theta(\log(1/\lambda))$ apart. Thus depending on the value of $\rho$, the new “Galton-Watson tree of high degree vertices” will either be supercritical or subcritical.

### v2

@@ Abstract
Scale-free configuration models are intimately connected to power law Galton-Watson trees. It is known that contact process epidemics can propagate on these trees and therefore these networks with arbitrarily small infection rate, and this continues to be true after uniformly immunising a small positive proportion of vertices. So, we instead immunise those with largest degree: above a threshold for the maximum permitted degree, we discover the epidemic with immunisation has survival probability similar to without, by duality corresponding to comparable metastable density. With maximal degree below a threshold on the same order, this survival probability is severely reduced or zero.
@@
John Fernley [^1] Corresponding author.,***HUN-REN Rényi Alfréd Matematikai Kutatóintézet, Reáltanoda utca 13-15, 1053 Budapest, Hungary. fernley@renyi.hu  Emmanuel JacobÉcole Normale Supérieure de Lyon, 46 allée d’Italie, 69007 Lyon, France. emmanuel.jacob@ens-lyon.fr
June 22, 2024
@@ 1 Introduction
On a locally finite graph $G=(V,E)$ modelling individuals and their interactions, the contact process is a Markov process in continuous time $(\xi_{t})_{t\geq 0}$ on $\{0,1\}^{V}$ with $\xi_{t}(v)=1$ (resp. 0) if the individual $x$ is infected (resp. healthy) at time $t$. The dynamics are determined by each infected vertex recovering at rate $1$ and each edge $x\sim y$ with for example $x$ infected and $y$ healthy transmitting the infection (to $y$) at fixed rate $\lambda>0$. On an infinite rooted graph $(G,o)$, we will typically start the infection from only the root infected and look at the survival probability of the infection $P_{\rm surv}(\lambda):=\operatorname{\mathbb{P}}(\forall t\geq 0,\xi_{t}\neq\mathbf{0})$. This survival probability is monotonic in $\lambda$, so that the survival probability is positive if $\lambda$ is larger than a critical parameter $\lambda_{c}=\lambda_{c}(G)\in[0,+\infty]$ and zero if $\lambda<\lambda_{c}$. Note the result [liggett1978] gives $\lambda_{c}(\mathbb{Z}^{d})\leq\nicefrac{{2}}{{d}}$, and we have $\lambda_{c}\leq\lambda_{c}(\mathbb{N})<+\infty$ on any infinite graph by comparison to oriented percolation [zbMATH02246274].
It is easy to prove that $\lambda_{c}\geq\nicefrac{{1}}{{d}}>0$ on any graph with degrees bounded by $d$, however it is totally possible to have $\lambda_{c}=0$ on graphs with unbounded degrees. We exactly know when this happens in the case of Galton-Watson trees since the works of [huang2018contact] and [bhamidi2019survival]. Consider $G$ a Galton-Watson tree with supercritical offspring distribution $X$ (namely $\operatorname{\mathbb{E}}X>1$), and conditioned to be infinite (otherwise $\lambda_{c}(G)=\infty$). Then $\lambda_{c}(G)$ is:
  • positive a.s. if $X$ has an exponential tail, i.e. there exists $c>0$ such that $\operatorname{\mathbb{E}}e^{cX}<+\infty$.
  • 0 a.s. otherwise.
In the latter case, for any positive $\lambda$, the quenched survival probability $P_{\rm surv}(\lambda,G):=\operatorname{\mathbb{P}}(\forall t\geq 0,\xi_{t}\neq\mathbf{0}|G)$ is thus a.s. positive on the event that $G$ is infinite, and in particular the annealed survival probability $P_{\rm surv}(\lambda)=\operatorname{\mathbb{E}}[P_{\rm surv}(\lambda,G)]$ is also positive.
We now specify the model and consider a Galton-Watson tree $G$ which:
  • is still supercritical: $\operatorname{\mathbb{E}}[X]>1$.
  • follows a power-law offspring distribution, or more precisely
    $\mathbb{P}(X\geq k)\sim c_{X}k^{2-\tau}$ (1)
    for some $c_{X}>0$ and power-law exponent $\tau>2$.
@@
Such a distribution has no finite exponential moment, so the (annealed) survival probability $P_{\rm surv}(\lambda)$ is positive for any $\lambda>0$. However this probability decays to 0 as $\lambda\to 0$, and as an adaptation of Proposition 1.4 in [mountford2013metastable], we can provide the following precise asymptotics for $P_{\rm surv}(\lambda)$:
@@
$P_{\rm surv}(\lambda)=\begin{cases}\Theta\Big{(}\lambda^{\frac{\tau-2}{3-\tau}}\Big{)},&\quad\text{when }\tau\leq 2.5,\\ \Theta\Big{(}\frac{\lambda^{2\tau-4}}{\log^{\tau-2}(\nicefrac{{1}}{{\lambda}})}\Big{)},&\quad\text{when }2.5<\tau\leq 3,\\ \Theta\Big{(}\frac{\lambda^{2\tau-4}}{\log^{2\tau-4}(\nicefrac{{1}}{{\lambda}})}\Big{)},&\quad\text{when }\tau>3,\end{cases}$ (2)
where $\Theta$ stands for the usual Landau notation.
The goal of this paper is to see to what extent this result is affected when we remove the full descendants of the highest degree vertices, and thus replace the offspring distribution $X$ by $X\mathbbm{1}_{X\leq\mathbbm{t}}$ for some truncation level $\mathbbm{t}=\mathbbm{t}(\lambda)$. Note that we immediately obtain extinction of the contact process when $\mathbbm{t}\leq\nicefrac{{1}}{{\lambda}}$, and so we will typically be interested in the regime $\mathbbm{t}=\mathbbm{t}(\lambda)>\nicefrac{{1}}{{\lambda}}$.
@@
Consider a truncation threshold $\mathbbm{t}=\mathbbm{t}(\lambda)\to+\infty$ as $\lambda\to 0$. Write $P^{\mathbbm{t}}_{\rm surv}(\lambda)$ for the survival probability of the contact process with infection parameter $\lambda>0$, starting from only the root infected and run on the Galton-Watson tree with truncated offspring distribution $X\mathbbm{1}_{X\leq\mathbbm{t}}$
@@
    $P^{\mathbbm{t}}_{\rm surv}(\lambda)=\begin{cases}O\left(e^{-\lambda^{-2+\delta}}\right),&\mathbbm{t}\leq\rho_{1}\lambda^{-2}\log\frac{1}{\lambda},\\ \Theta\left(\frac{\lambda^{2\tau-4}}{\log^{\tau-2}\left(\nicefrac{{1}}{{\lambda}}\right)}\right),&\mathbbm{t}\geq\rho_{2}\lambda^{-2}\log\frac{1}{\lambda}.\end{cases}$
@@
    $P^{\mathbbm{t}}_{\rm surv}(\lambda)=\begin{cases}O\left(\lambda^{C}\right),&\mathbbm{t}\leq\rho_{1}\lambda^{-2}\log^{2}\frac{1}{\lambda},\\ \Theta\left(\frac{\lambda^{2\tau-4}}{\log^{2\tau-4}\left(\nicefrac{{1}}{{\lambda}}\right)}\right),&\mathbbm{t}\geq\rho_{2}\lambda^{-2}\log^{2}\frac{1}{\lambda}.\end{cases}$
@@
These are asymptotic results as $\lambda\to 0$ and the truncation threshold has $\mathbbm{t}=\mathbbm{t}(\lambda)\to+\infty$. In particular, the truncated degree distribution $X\mathbbm{1}_{X\leq\mathbbm{t}}$ will have first moment larger than 1 and the truncated Galton-Watson tree will be supercritical, as is necessary for the contact process to have any chance to survive. More generally and even when not explicitly stated, we will generally work on the assumption that $\lambda>0$ is sufficiently small.
@@
In the case $\tau<2.5$, if we also choose $\mathbbm{t}=\rho\lambda^{-\frac{1}{3-\tau}}$, we thus obtain a phase transition in $\rho$ where:
  • For $\rho$ smaller than the critical value $\rho_{c}$, the contact process gets extinct almost surely
  • For $\rho$ larger than the critical value $\rho_{c}$, the truncation does not have any appreciable effect on the survival probability, which is at most changed by a multiplicative constant when $\lambda\to 0$ as compared to $P_{\rm surv}(\lambda)$.
@@
In the cases where $\tau\geq 2.5$, we believe a similar phase transition occurs, from almost sure extinction of the contact process, to a survival probability which is essentially the same as without truncation, and in line with (2). However, our results are not as complete in those cases. First, we do not reveal a single critical value for $\rho$, but rather have to suppose $\rho$ large to get our large $\mathbbm{t}$ estimates, and $\rho$ small to obtain our small $\mathbbm{t}$ upper bounds. Second, if $\tau>2.5$, we cannot even prove extinction of the process for small $\rho$, and instead we only get that the survival probability is significantly decreased. This is due to the difficulty of proving extinction of the contact process under some very unlikely but still possible events, like the truncated tree having many vertices with maximal degree adjacent to the root. Using our current techniques, we still can prove almost sure extinction of the infection, but only if we truncate at a much smaller degree $\mathbbm{t}=O(\lambda^{-2})$.
@@
In the physics literature, these questions have been considered by [pastor2002epidemic, pastor2002immunization, deszo2003halting] using mean-field heuristics. However, as previously observed by [chatterjee2009contact] for the problem without immunisation, these heuristics struggle to accurately account for the effect of vertices of large degree. Targeted immunisation of the large degree vertices is therefore where they are most difficult to justify.
This has led to estimates [deszo2003halting, Equation 2] for general networks with $2<\tau\leq 3$ or [pastor2002immunization, Equation 28] for a preferential atttachment network with $\tau=3$, both of the form
$\mathbbm{t}=e^{\Theta\left(\nicefrac{{1}}{{\lambda}}\right)}$
and so starkly disagreeing with the polynomial order $\mathbbm{t}$ of our main theorem.
@@
Consider immunization instead of truncation, and write $\widetilde{P}^{\mathbbm{t}}_{\rm surv}(\lambda)$ for the survival probability of the contact process with infection parameter $\lambda>0$, starting from only the root infected, on the Galton-Watson tree with offspring distribution $X$, but where every vertex with offspring [^2] We could also immunize every vertex with degree larger than $\mathbbm{t}+1$, instead of every vertex with offspring larger than $\mathbbm{t}$. This would make a difference only for the root (if it has degree exactly $\mathbbm{t}+1$), but would not change our results. larger than $\mathbbm{t}$ is immunized. By definition the immunized vertices cannot get infected or transmit the infection, so these vertices and their offspring can just be removed from the tree. Note this is different from truncation, where only the offspring of the truncated vertices were removed. In principle the contact process could have a different behaviour on the truncated tree, as a truncated vertex can still become infected and reinfect its parent. However, truncation or immunization make no difference when looking at our results.
Claim 1.6.
@@
Remark 1.7.
Let $X$ follow the common offspring distribution. As discussed previously, the truncated tree is simply a Galton-Watson tree with offspring distribution $X\mathbbm{1}_{X\leq\mathbbm{t}}$. The tree obtained after removal of the immunized vertices is also a Galton-Watson tree, but with a more complicated offspring distribution, which is the main reason for us to consider truncation afterwards. Indeed, let $Y$ have the law of $X$ conditionally on $X\leq\mathbbm{t}$. Then each vertex has probability $p=\operatorname{\mathbb{P}}(X\leq\mathbbm{t})$ to be a small vertex (with offspring $\leq\mathbbm{t}$), and conditionally on being a small vertex, has offspring distribution $Y$, each child being again a small vertex with probability $p$. Hence, if we condition on the root to be a small vertex, then the tree obtained after removal of the big vertices is a Galton-Watson tree with mixed binomial offspring distribution $\operatorname{Bin}\left(Y,p\right).$
@@ 2 Overview
2.1 The SIR Galton-Watson tree
The SIR process is simply the contact process with permanent recoveries, which is a stochastic lower bound [liggett1999stochastic, Section I.1] . On a tree with root $o$ and initial infection $\xi_{0}=\mathbbm{1}_{o}$, this simplifies the analysis in that infection can only move away from the root. In fact, simply give each vertex an $\operatorname{Exp}(1)$ recovery time and each vertex but the root an $\operatorname{Exp}(\lambda)$ infection time from its parent, and then remove every edge for which the infection time of the child is larger than the recovery time of the parent. The connected component of $o$ is then the subtree of vertices which receive the infection.
Hence, putting the above construction in a Galton-Watson tree produces a Galton-Watson subtree which is the easiest lower bound for our problem.
Definition 2.1.
Draw $D$ from some offspring distribution, and $I\sim\operatorname{Exp}(1)$. Then the offspring distribution of the SIR Galton-Watson tree is $Y$ with distribution
$Y|D,I\sim\operatorname{Bin}\left(D,1-e^{-\lambda I}\right).$
@@ 2.2 Local survival by reinfection from your neighbours, and survival strategies
On the star graph containing a central vertex and $k=k(\lambda)$ leaves, with initially only the center infected, the center will typically not be permanently recovered before time
@@
That is to say, they will typically see the infection at some time later than such a $T$.
This simple result dates back to [pemantle1992contact, Theorem 4.1] and has been rewritten or detailed several times since, see e.g. [berger2005starsurvival, Lemma 5.2] or [nguyen2022subcritical, Lemma 3.4]. See also [periodic2020, Lemma 4] for a similar upper bound for the expected extinction time of the epidemics. We will use a version of this result stated for the total time the center is infected, see Lemma A.1, which is actually a refinement of [nguyen2022subcritical, Lemma 3.4] of independent interest.
On a larger graph, $T(\lambda,\deg(x))$ can still be interpreted heuristically as the time that the infection can survive “locally” at a vertex $x$, using only direct reinfections from the neighbours of $x$. On the truncated tree, the maximal such value is of course $T(\lambda,\mathbbm{t}+1)$, which differs a lot in the different regimes considered in our main theorem. Writing $D$ for the offspring distribution of the tree environment:
@@
    Case $\tau\leq 2.5$: Consider $\mathbbm{t}=\rho\lambda^{-\frac{1}{3-\tau}}$ for some constant $\rho>0$. In the case $\tau<2.5$ we have $T(\lambda,\mathbbm{t}+1)\to 1$ as $\lambda\to 0$, which is an indication that the reinfections barely have any effect on the spread of the epidemics. We thus expect to survive as in an SIR process, namely if and only if we have $\lambda\operatorname{\mathbb{E}}[D]>1$. When $\tau=2.5$ we now have $T(\lambda,\mathbbm{t}+1)=e^{\Theta(\rho)}$ of constant order. We still expect survival by the SIR lower bound if $\lambda\operatorname{\mathbb{E}}[D]>1$, while we expect extinction if we have both $\lambda\operatorname{\mathbb{E}}[D]<1$ and $\rho$ sufficiently small.
@@
    Case $2.5<\tau\leq 3$. Consider $\mathbbm{t}=\rho\lambda^{-2}\log\nicefrac{{1}}{{\lambda}}$. We now have $T(\lambda,\mathbbm{t}+1)=\lambda^{-\Theta(\rho)}$. When $\rho$ is large, this time is larger than $\lambda^{-1}$ and thus sufficient to typically infect any given neighbour. We then ensure survival of the infection by observing that in that case the high-degree vertices (with degree of order $\mathbbm{t}$) always form a supercritical Galton-Watson process on which the infection can spread.
    On the contrary, when $\rho$ is small, we infect any given neighbour with probability only $\lambda^{1-\Theta(\rho)}$, which is not much larger than $\lambda$. In this phase we always have $\lambda\operatorname{\mathbb{E}}[D]\to 0$ as well as $\lambda^{1-\Theta(\rho)}\operatorname{\mathbb{E}}[D]\to 0$ when $\rho$ is small. Therefore the strategy of maintaining the infection at each vertex $x$ for time $T(\lambda,\deg(x))$ before passing the infection to its children will not succeed, which is an indication that the contact process might die out.
@@
    $T(\lambda,\mathbbm{t}+1)=\lambda^{-\Theta(\rho)\log(\nicefrac{{1}}{{\lambda}})},$
    a time which is sufficient to infect not only the direct neighbours, but also all nearby vertices up to distance $\Theta(\rho\log(\nicefrac{{1}}{{\lambda}}))$. It is thus natural to consider that two high-degree vertices are “connected” if they are separated by less than $\Theta(\rho\log(\nicefrac{{1}}{{\lambda}}))$ small-degree vertices, hence obtaining a new “Galton-Watson tree of high degree vertices”.
    As $\tau>3$ the high-degree vertices are more rare in the original tree, and actually typically at distance $\Theta(\log(\nicefrac{{1}}{{\lambda}}))$ apart. Thus depending on the value of $\rho$, the new “Galton-Watson tree of high degree vertices” will either be supercritical or subcritical.